# OTTER: Improving Zero-Shot Classification via Optimal Transport

### v1

@@ Abstract
Popular zero-shot models suffer due to artifacts inherited from pretraining. A particularly detrimental artifact, caused by unbalanced web-scale pretraining data, is mismatched label distribution. Existing approaches that seek to repair the label distribution are not suitable in zero-shot settings, as they have incompatible requirements such as access to labeled downstream task data or knowledge of the true label balance in the pretraining distribution. We sidestep these challenges and introduce a simple and lightweight approach to adjust pretrained model predictions via optimal transport. Our technique requires only an estimate of the label distribution of a downstream task. Theoretically, we characterize the improvement produced by our procedure under certain mild conditions and provide bounds on the error caused by misspecification. Empirically, we validate our method in a wide array of zero-shot image and text classification tasks, improving accuracy by 4.8% and 15.9% on average, and beating baselines like Prior Matching—often by significant margins—in 17 out of 21 datasets.
@@ 1 Introduction
Zero-shot models show promise for a variety of applications where training data is challenging to collect or label. After pretraining a model on data-rich tasks, the model is presented with an unseen, downstream task dataset at inference time, eliminating the need for data annotation and model training. Zero-shot learning has been successfully applied to data- or compute- limited tasks ranging from image retrieval [49, 51, 55, 54], to video event detection [15, 10, 39, 57] and spoken language understanding [69, 25].
Unfortunately, zero-shot models inherit bias from their large pretraining datasets [19, 62, 2]—which are typically composed of massive web-crawled data. In particular, zero-shot classification is strongly biased by the label distribution of the pretraining task. When the label distribution of the downstream task differs from pretraining, the performance of zero-shot classifiers suffers greatly. For example, Figure 1 illustrates the effects of mismatched distributions on a pet image classification task. CLIP (RN50) and CLIP (ViT-B/16) produce biased predictions on Abyssinian and Persian, owing to the distribution of images and words in their pretraining data. However, the pet distribution is uniformly distributed in this task. This adversely affects their accuracy in such tasks. Meanwhile, datasets with a large number of classes, such as ImageNet, may contain both extremely common and very rare classes, resulting in an outsized probability that a zero-shot model will predict some classes over others. As a result, even large models intended for use in zero-shot settings, such as CLIP [48], naturally have a label distribution mismatch between pretraining data and downstream tasks.
Figure 1: Label distribution mismatch example in zero-shot classification. In Oxford-IIIT-Pet dataset, the labels are uniformly distributed, while the zero-shot models exhibit biased predictions toward some clasess. This bias is influenced by the distribution of labels in the pretraining task.
Existing methods that seek to address label distribution mismatch include fine-tuning and label shift adaptation [33, 6, 22]. The latter reweights the classifier scores with the ratio of the estimated label distribution in the target distribution and the true label distribution in the source distribution. However, these approaches make strong assumptions: to fine-tune a model, we must obtain a labeled fine-tuning dataset of adequate size, then possess the time and compute to train the model further. To perform label shift adaptation, we must know the true label distribution of the pretraining distribution—difficult to impossible for real-world tasks.
Can we deal with label distribution mismatch without additional training or access to ground-truth downstream task information? One cause for optimism is the observation that zero-shot models still give relatively high prediction probabilities for correct classes, though classes common in pretraining tend to have relatively inflated scores overall. Intuitively, the model has already learned to identify examples of its downstream classes (and so does not require further training) and is already impacted by the pretraining label distribution (and so does not need access to the ground-truth pretraining label distribution). Instead, the model’s prediction probabilities must be adjusted based on an estimated downstream label distribution specification.
@@
Theoretically, we show that optimal transport given the true label distribution of the downstream can recover the Bayes-optimal classifier under mild conditions. Additionally, we provide error bounds on our adaptation method for misspecification. We provide synthetic experiments validating our theoretical claims. In real-world experiments, we validate our method on a wide variety of image and text classification tasks, showing 4.8% and 15.5% accuracy improvement on average in image and text zero-shot classification tasks, respectively. Finally, we evaluate our method in few-shot adaptation scenarios, where OTTER provides further improvements when combined with linear probing. Our contributions include:
@@
  • Theoretical results showing the effectiveness of our method, including the ability to recover the Bayes-optimal classifier and a sensitivity analysis with respect to the label distribution specification estimations,
  • Extensive empirical results on zero-shot classification for text and image datasets, showing accuracy improvements of up to 25% and
  • Experimental results demonstrating the applicability of OTTER to few-shot settings, showing accuracy improvements of up to 15%, even with noisy label distribution specification.
@@ 2.1 Background
Zero-shot classification, popularized by models such as CLIP [48], is a powerful paradigm that enables prediction on downstream tasks without additional fine-tuning. Image, language, and multimodal models have been increasingly employed for zero-shot prediction [66, 34]. These models undergo extensive pretraining on massive datasets with label spaces different from those of downstream applications.
@@
Optimal Transport (OT) is a framework for matching two probability distributions [47, 56]. We predominantly consider optimal transport between empirical discrete measures. Suppose that we are given points $x_{1},x_{2},\ldots,x_{n}\in\mathcal{X}$ and $y_{1},\ldots,y_{K}\in\mathcal{Y}$, a source measure $\mu$ defined by $\mu=\sum_{i=1}^{n}w_{i}\delta_{x_{i}}$, and a target measure given by $\nu=\sum_{j=1}^{K}p_{j}\delta_{j}$, where $w_{i},p_{j}$ are positive values such that $\sum_{i=1}^{n}w_{i}=1$, $\sum_{j=1}^{K}p_{j}=1$. Suppose also that $\delta_{x}$ is a Dirac delta function at $x$, i.e. $\delta_{x}(x^{\prime})=1$ if $x=x^{\prime}$, and $\delta_{x}(x^{\prime})=0$ otherwise. Given a cost matrix $C\in\mathbb{R}^{n\times K}$, the Monge-Kantorovich formulation of optimal transport is to find a minimal cost transport plan $\pi$ such that
@@
Distribution shift refers to the discrepancy between a source distribution $P_{s}$ on which the model is trained, and a target distribution $P_{t}$ on which the model is deployed. Distribution shift often degrades trained model performance on target tasks. Label shift is a specific type of distribution shift such that $P_{s}(Y)\neq P_{t}(Y)$ and the data generation process is fixed — in other words, the conditional distributions of the inputs are the same: $P_{s}(X|Y)=P_{t}(X|Y)$. Techniques such as importance sampling [33, 6, 22], recalibration [3] and domain adaptation [59] are commonly used to mitigate the effects of label shift. Unfortunately, these methods assume access to source distribution data, whereas zero-shot models’ pretraining data is often inaccessible (often proprietary or blocked for privacy reasons). Thus, adapting zero-shot models to new label distributions poses new challenges unmet by these pre-existing methods.
@@ 2.2 Problem Formulation
Let $X=\{x_{1},x_{2},\ldots,x_{n}\}$ be an inference dataset with $x_{i}\in\mathcal{X}$. Furthermore, let $Y=\{y_{1},y_{2},\ldots,y_{n}\}$ be the true labels of the $K$-class classification dataset, such that $y_{i}\in\mathcal{Y}=[K]$, are sampled according to the downstream label distribution $\nu=(p_{1},p_{2},\ldots,p_{K})$.
Let $s_{\theta}(x,j):=P_{\theta}(Y=j|X=x)$ be a pretrained classification model constrained to the downstream label space. During pretraining, $s_{\theta}$ has been biased to the source label distribution $\nu^{s}$. We wish to offset such label distribution bias with a label distribution specification $\hat{\nu}$ in target distribution. $\hat{\nu}$ is expected to be closer to the true label distribution of the downstream task. Given a label distribution specification, our goal is to rebalance predictions so that the predicted label distribution follows the label distribution specification.
@@ 3 Proposed Framework
3:Run optimal transport and obtain transport plan $\pi$ s.t.
$\pi=\operatorname*{arg\,min}_{\gamma\in\Pi(\mu,\nu)}\langle\gamma,C\rangle.$
4:Get modified classification outputs
$\hat{y}_{i}=\arg\max_{j\in[K]}\pi_{i,j}.$
@@
We propose OTTER (Optimal TransporT adaptER), an optimal transport-based label distribution adaptation approach. Our goal is to have the $n$ input data points allocated to $K$ classes to match a given label distribution $\hat{\nu}$, where $\sum_{j=1}^{K}\hat{\nu}_{j}=1,\hat{\nu}_{j}\geq 0$. Specifically, we want to classify $n\hat{\nu}_{1}$ points as class 1, $n\hat{\nu}_{2}$ points as class 2, and so on. However, there are many such allocations, and it is not a priori clear which one should be selected. We propose formulating an optimal transport problem that selects the allocation minimizing a particular cost:
@@
We wish to use Algorithm 1 for zero-shot classification, given the pretrained model $s_{\theta}$. To do so, we must select a cost function and produce $C_{ij}$. An ideal choice of such a function is $C_{ij}=-\log_{t}P_{t}(Y=j|X=i)$ such that optimal transport minimizes the negative log posterior under constraints. However, the target distribution $P_{t}$ is unknown. Instead, we replace the posterior with the classifier scores $s_{\theta}(x_{i},j)$. We highlight that this choice of cost matrix is an natural extension of zero-shot classification under the label distribution constraint. We theoretically prove this claim in the following section. Next, we provide a toy example that shows how OTTER can improve zero-shot accuracy.
@@
To illustrate the benefit of OTTER, consider the following example for binary classification. We have two datapoints, $X=\{x_{1},x_{2}\}$ with $Y=\{1,2\}$, and true label distribution $\nu=(\frac{1}{2},\frac{1}{2})$. Suppose that the zero-shot model’s prediction scores are $s_{1}=(0.4,0.6)$ and $s_{2}=(0.1,0.9)$.
@@
$S=\begin{bmatrix}0.4&0.6\\ 0.1&0.9\end{bmatrix},C=\begin{bmatrix}-\log 0.4&-\log 0.6\\ -\log 0.1&-\log 0.9\end{bmatrix}$
@@ 4 Theoretical Results
In practical scenarios, label distribution specifications are frequently subject to noise, and prediction probabilities may not be well-calibrated. To understand the impact of these factors, we examine how errors in label distribution specification and calibration influence the transport plan. Our comprehensive theoretical analysis yields following findings: (a) OTTER can recover the Bayes-optimal classifier in the label shift setting, (b) for a noisy cost matrix with the noisy label distribution specification setup, the suboptimality can be bounded by the deviation of cost matrix and label distribution.
@@
Suppose $\nu^{ZS}_{j}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}[\hat{y}^{ZS}_{i}=j]$, where $\hat{y}^{ZS}_{i}=\arg\max_{j^{\prime}\in[K]}P_{\theta}(Y=j^{\prime}|X=x_{i})$. Then, given $C_{ij}=-\log P_{\theta}(Y=j|X=x_{i})$,
@@
Assuming there are no ties in scores, i.e. $P_{\theta}(Y=j|X=x_{i})\neq P_{\theta}(Y=j^{\prime}|X=x_{i}),\text{ for all }j\neq j^{\prime}$, the predictions are equivalent, i.e. $\hat{y}^{OT}_{i}=\hat{y}^{ZS}_{i}$ for all $i\in[n]$.
This theorem has the following implications. First, it suggests that the predictions will remain unchanged if we set $\hat{\nu}=\nu^{ZS}$. Secondly, Bayes-optimal classifiers can be derived through optimal transport, using a (true) cost matrix defined as $C^{*}_{ij}=-\log P_{t}(Y=j|X=x_{i})$, coupled with the true label distribution $\nu^{*}$.
Our analysis begins with the label shift setup, which is a commonly-studied type of distribution shift—as well as a prominent issue when applying zero-shot models. We demonstrate that when the label distribution is correctly specified, optimal transport preserves the Bayes-optimal classifier predictions under label shift. Next, we consider general perturbations in label distribution and cost matrix as well as their impact on the resulting solutions.
@@ 4.1 Label Shift Invariance
$\tilde{s}_{\theta}(x,j)$ $=s_{\theta}(x,j)\cdot\frac{P_{t}(Y=j)}{P_{s}(Y=j)}$
$\propto P_{s}(Y=j,X=x)\cdot\frac{P_{t}(Y=j)}{P_{s}(Y=j)}$
$=P_{s}(X=x|Y=j)\cdot P_{s}(Y=j)\cdot\frac{P_{t}(Y=j)}{P_{s}(Y=j)}$
$=P_{t}(X=x|Y=j)\cdot P_{t}(Y=j)$
$\propto P_{t}(Y=j|X=x).$
@@
Denote the Bayes optimal predictions in the target distribution as $\hat{y}^{*}_{i}=\arg\max_{j\in[K]}\log P_{t}(Y=j|X=x_{i})$. Then OTTER predictions $\textsc{OTTER}(\mathbf{X},P_{t}(Y),C)$ are the same as Bayes optimal predictions $\hat{y}^{*}$.
Theorem 4.2 implies that OTTER recovers a Bayes classifier in the target distribution without access to the source distribution, given the target label distribution and a model well-calibrated for the source dataset.

### v2

@@ Abstract
Popular zero-shot models suffer due to artifacts inherited from pretraining. One particularly detrimental issue, caused by unbalanced web-scale pretraining data, is mismatched label distribution. Existing approaches that seek to repair the label distribution are not suitable in zero-shot settings, as they have mismatching requirements, such as needing access to labeled downstream task data or knowledge of the true label balance in the pretraining distribution. We sidestep these challenges and introduce a simple and lightweight approach to adjust pretrained model predictions via optimal transport. Our technique requires only an estimate of the label distribution of a downstream task. Theoretically, we characterize the improvement produced by our procedure under certain mild conditions and provide bounds on the error caused by misspecification. Empirically, we validate our method in a wide array of zero-shot image and text classification tasks, improving accuracy by 4.8% and 15.9% on average, and beating baselines like prior matching—often by significant margins—in 17 out of 21 datasets.
@@ 1 Introduction
Figure 1: Label distribution mismatch example in zero-shot classification. In the Oxford-IIIT-Pet dataset, the ground-truth labels are uniformly distributed, while zero-shot models exhibit biased predictions toward certain classes. This bias is influenced by the distribution of labels in the pretraining task.
Zero-shot models are popular but struggle with biases inherited from their large pretraining datasets [bias_1, bias_2, bias_inherit_1]. In particular, zero-shot classification is strongly biased by the label distribution of the pretraining task. When the label distribution of the downstream task differs from pretraining, the performance of zero-shot classifiers suffers greatly. For example, Figure 1 illustrates the effects of mismatched distributions on a pet image classification task. Two CLIP models (RN50, and ViT-B/16) produce biased predictions on the Abyssinian and Persian classes. Furthermore, datasets with a large number of classes, such as ImageNet, may contain both extremely common and very rare classes, resulting in an outsized probability that a zero-shot model will predict some classes over others. As a result, even large models intended for use in zero-shot settings, such as CLIP [CLIP_base], naturally have a label distribution mismatch between pretraining data and downstream tasks.
Existing methods that seek to address label distribution make strong assumptions or have expensive requirements. For example, to fine-tune a model, we must obtain a labeled fine-tuning dataset of adequate size, then obtain the time and compute to further train the model. To perform label shift adaptation techniques, we must know the true label distribution of the pretraining distribution—difficult to impossible for real-world tasks.
Can we deal with label distribution mismatch without additional training or access to ground-truth downstream task information? While seemingly challenging, one cause for optimism is the observation that zero-shot models still give relatively high prediction probabilities for correct classes, though classes common in pretraining tend to have relatively inflated scores overall. Intuitively, the model has already learned to identify examples of its downstream classes (and so does not require further training) and is already impacted by the pretraining label distribution (and so does not need access to the ground-truth pretraining label distribution). Instead, the model’s prediction probabilities must be adjusted based on an estimated downstream label distribution specification.
@@
Theoretically, we show that optimal transport given the true label distribution of the downstream can recover the Bayes-optimal classifier under mild conditions. Additionally, we provide error bounds on our adaptation method for misspecification. We provide synthetic experiments validating our theoretical claims. In real-world data settings, we validate our method on a wide variety of image and text classification tasks, showing 4.8% and 15.5% accuracy improvement on average in image and text zero-shot classification tasks, respectively. Finally, we evaluate our method in few-shot adaptation scenarios, where OTTER provides further improvements when combined with linear probing. Our contributions include:
@@
  • Theoretical results showing the effectiveness of our method, including the ability to recover the Bayes-optimal classifier and a sensitivity analysis with respect to the label distribution specification estimation,
  • Extensive empirical results on zero-shot classification for text and image datasets, showing accuracy improvements of up to 25%,
  • Experimental results demonstrating the applicability of OTTER to few-shot settings, showing accuracy improvements of up to 15%, even with noisy label distribution specifications,
  • Extensions of OTTER to leverage label hierarchy information or relax the batched prediction requirement and
  • Application to LLM selection bias mitigation.
@@ 2.1 Background
We briefly describe zero-shot models, the technical tool we use (optimal transport), along with other techniques that seek to address shifts. We have extended related work, including in-depth characterizations and comparisons with related methods, in Appendix LABEL:sec:related_works.
@@
Zero-shot classification, popularized by models such as CLIP [CLIP_base], is a powerful paradigm that enables prediction on downstream tasks without additional fine-tuning. Image, language, and multimodal models have been increasingly employed for zero-shot prediction [wei2021finetuned, liu2023visual]. These models undergo extensive pretraining on massive datasets with concept and label spaces that may be very different from those of downstream applications.
@@
Optimal Transport (OT) is a framework for matching two probability distributions [peyre2019computational, santambrogio2015optimal]. We predominantly consider optimal transport between empirical discrete measures. Suppose that we are given points $x_{1},x_{2},\ldots,x_{n}\in\mathcal{X}$ and $y_{1},\ldots,y_{K}\in\mathcal{Y}$, a source measure $\mu$ defined by $\mu=\sum_{i=1}^{n}w_{i}\delta_{x_{i}}$, and a target measure given by $\nu=\sum_{j=1}^{K}p_{j}\delta_{j}$, where $w_{i},p_{j}$ are positive values such that $\sum_{i=1}^{n}w_{i}=1$, $\sum_{j=1}^{K}p_{j}=1$. Suppose also that $\delta_{x}$ is a Dirac delta function at $x$, i.e. $\delta_{x}(x^{\prime})=1$ if $x^{\prime}=x$, and $\delta_{x}(x^{\prime})=0$ otherwise. Given a cost matrix $C\in\mathbb{R}^{n\times K}$, the Monge-Kantorovich formulation of optimal transport is to find a minimal cost transport plan $\pi$ such that
@@
Distribution shift refers to the discrepancy between a source distribution $P_{s}$ on which the model is trained, and a target distribution $P_{t}$ on which the model is deployed. Distribution shift often degrades trained model performance on target tasks. Label shift is a specific type of distribution shift such that $P_{s}(Y)\neq P_{t}(Y)$ and the data generation process is fixed — in other words, the conditional distributions of the inputs are the same: $P_{s}(X|Y)=P_{t}(X|Y)$. Techniques such as importance sampling [lipton2018detectingbbse, azizzadenesheli2019regularizedrlls, garg2020unifiedmlls], recalibration [alexandari2020maximum] and domain adaptation [tachet2020domain] are commonly used to mitigate the effects of label shift. Unfortunately, these methods assume access to source distribution data, whereas zero-shot models’ pretraining data is inaccessible (often proprietary or blocked for privacy reasons). Thus, adapting zero-shot models to new label distributions poses challenges unmet by these pre-existing methods.
@@ 2.2 Problem Formulation
Let $\mathbf{X}=\{x_{1},x_{2},\ldots,x_{n}\}$ be an inference dataset with $x_{i}\in\mathcal{X}$. Furthermore, let $\mathbf{Y}=\{y_{1},y_{2},\ldots,y_{n}\}$ be the true labels of the $K$-class classification dataset, such that $y_{i}\in\mathcal{Y}=[K]$, are sampled according to the downstream label distribution $\nu=(p_{1},p_{2},\ldots,p_{K})$.
Let $s_{\theta}(x,j):=P_{\theta}(Y=j|X=x)$ be a pretrained classification model constrained to the downstream label space. During pretraining, $s_{\theta}$ has been biased to the source label distribution $\nu^{s}$. We wish to offset such label distribution bias with a label distribution specification $\hat{\nu}$ for the target distribution. $\hat{\nu}$ is expected to be closer to the true label distribution of the downstream task. Given a label distribution specification, our goal is to rebalance predictions so that the predicted label distribution follows the label distribution specification.
@@ 3 Proposed Framework
3:Run optimal transport and obtain transport plan $\pi$ s.t. $\pi=\operatorname*{arg\,min}_{\gamma\in\Pi(\mu,\nu)}\langle\gamma,C\rangle$.
4:Get modified classification outputs $\hat{y}_{i}=\arg\max_{j\in[K]}\pi_{i,j}$.
@@
We propose OTTER (Optimal TransporT adaptER), an optimal transport-based label distribution adaptation approach. Our goal is to have the $n$ input data points allocated to $K$ classes match a given label distribution $\hat{\nu}$, where $\sum_{j=1}^{K}\hat{\nu}_{j}=1,\hat{\nu}_{j}\geq 0$. Specifically, we want to classify $n\hat{\nu}_{1}$ points as the first class, $n\hat{\nu}_{2}$ points as the second, and so on. However, there are many such allocations, and it is not a priori clear which one should be selected. We propose formulating an optimal transport problem that selects the allocation minimizing a particular cost:
@@
We wish to use Algorithm 1 for zero-shot classification given the pretrained model $s_{\theta}$. To do so, we must select a cost function and produce $C_{ij}$. An ideal choice of such a function is $C_{ij}=-\log P_{t}(Y=j|X=i)$ such that optimal transport minimizes the negative log posterior under constraints. However, the target distribution $P_{t}$ is unknown. Instead, we replace the posterior with the classifier scores $s_{\theta}(x_{i},j)$. We highlight that this choice of cost matrix is an natural extension of zero-shot classification under the label distribution constraint. We prove this claim in the next section. First, we show a toy example that shows how OTTER improves zero-shot accuracy.
@@
To illustrate the benefits of OTTER, consider the following example for binary classification. We have two data points, $X=\{x_{1},x_{2}\}$ with $Y=\{1,2\}$, and true label distribution $\nu=(\frac{1}{2},\frac{1}{2})$. Suppose that the zero-shot model’s prediction scores are $s_{1}=(0.4,0.6)$ and $s_{2}=(0.1,0.9)$.
@@
$S=\begin{bmatrix}0.4&0.6\\ 0.1&0.9\end{bmatrix},C=\begin{bmatrix}-\log 0.4&-\log 0.6\\ -\log 0.1&-\log 0.9\end{bmatrix},$
@@
Extension to Online Predictions
[…]
$P_{\theta,r}(Y=j|X=x)=\cfrac{r_{j}P_{\theta}(Y=j|X=x)}{\sum_{j^{\prime}=1}^{K}r_{j^{\prime}}P_{\theta}(Y=j^{\prime}|X=x)}.$
[…]
@@ 4 Theoretical Results
[…]
@@
Let $\nu^{ZS}_{j}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}[\hat{y}^{ZS}_{i}=j]$, where $\hat{y}^{ZS}_{i}=\arg\max_{j^{\prime}\in[K]}P_{\theta}(Y=j^{\prime}|X=x_{i})$. Then, given $C_{ij}=-\log P_{\theta}(Y=j|X=x_{i})$,
@@
Assuming there are no ties in scores, i.e. $P_{\theta}(Y=j|X=x_{i})\neq P_{\theta}(Y=j^{\prime}|X=x_{i}),\text{ for all }j\neq j^{\prime}$, the OTTER predictions are equivalent zero-shot predictions, i.e. $\hat{y}^{OT}_{i}=\hat{y}^{ZS}_{i}$ for all $i\in[n]$.
This theorem has the following implications. First, it suggests that the predictions will remain unchanged if we set $\hat{\nu}=\nu^{ZS}$. Second, Bayes-optimal classifiers can be derived through optimal transport, using a (true) cost matrix defined as $C^{*}_{ij}=-\log P_{t}(Y=j|X=x_{i})$, coupled with the true label distribution $\nu^{*}$.
Our analysis begins with the label shift setup, which is a commonly-studied type of distribution shift—as well as a prominent issue when applying zero-shot models. We demonstrate that when the label distribution is correctly specified, optimal transport preserves Bayes-optimal classifier predictions under label shift. Next, we consider general perturbations in label distribution and cost matrix as well as their impact on the resulting solutions.
@@ 4.1 Label Shift Invariance
$\tilde{s}_{\theta}(x,j)$ $=s_{\theta}(x,j)\cdot\frac{P_{t}(Y=j)}{P_{s}(Y=j)}\propto P_{t}(X=x|Y=j)\cdot P_{t}(Y=j)\propto P_{t}(Y=j|X=x).$
@@
Denote the Bayes optimal predictions in the target distribution as $\hat{y}^{*}_{i}=\arg\max_{j\in[K]}\log P_{t}(Y=j|X=x_{i})$. Then OTTER predictions $\hat{y}=\textsc{OTTER}(\mathbf{X},\nu^{*},C)$ are the same as Bayes optimal predictions $\hat{y}^{*}$.
That is, OTTER recovers a Bayes classifier in the target distribution without access to the source distribution, given the target distribution and a well-calibrated model for the source dataset.